Let f(x)=x-7/x+5 and g(x)=x+3/x-4
f(x)=g(x)
I am stumped and maybe if i got the answer it could help me learn the process.

Answer by Alan3354(69443)

Let f(x)=x-7/x+5 and g(x)=x+3/x-4
f(x)=g(x)
---------------
x-7/x+5 = x+3/x-4
(x-7)*(x-4) = (x+5)*(x+3)
x^2 - 11x + 28 = x^2 + 8x + 15
-11x + 28 = 8x + 15
-19x = -13
x = 13/19